OpenStudy (anonymous):

Home plate on a baseball field has three right angles and two congruent angles. Find the measure of the two congruent angles

OpenStudy (anonymous):

Hmm, I always thought it had 2 right angles, but I suppose the point that points to the catcher is a right angle It looks like this http://beaconathletics.com/store/images/products/301-675-859.jpg If you know the measure of the interior angles of a n-gon (in our case, it is 5), you can do 3*90 + 2x = measure of interior angles of the pentagon does that help?

OpenStudy (anonymous):

Yeah that kind of helps would I just use that equation?

OpenStudy (anonymous):

I looked online, and the formula is ((n-2)*180) = sum of interior angles http://www.mathopenref.com/polygoninteriorangles.html so the sum of the home plate should be 3(90) + 2(x) = ((5-2)*180)) I'll let you solve the rest!

Directrix (directrix):

Home Plate is a convex pentagon. The pentagon has an interior angle sum of 540. 540 - 3(90) = 540 - 270 = 270 for the sum of the remaining two congruent angles. 270/2 = 135 degrees for those two.